# Stochastic Online Metric Matching: Adversarial is no Harder than Stochastic

### v1

@@ Abstract
We study the stochastic online metric matching problem. In this problem, $m$ servers and $n$ requests are located in a metric space, where all servers are available upfront and requests arrive one at a time. In particular, servers are adversarially chosen, and requests are independently drawn from a known distribution. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to a free server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
In this paper, we show that the problem can be reduced to a more accessible setting where both servers and requests are drawn from the same distribution by incurring a moderate cost. Combining our reduction with previous techniques, for $[0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regrets in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of [GGPW19] for balanced markets in various regimes, and provide the first positive results for unbalanced markets.
@@ 1 Introduction
Online matching forms a central topic of online algorithms and has received extensive attention due to the surge of online matching markets in recent years. Some of its prominent applications include online advertising, job search, and dating, where the constant growth of online advertising platforms has particularly spurred numerous research from both algorithmic and economic perspectives. A seminal paper [KVV90] proposes the online bipartite matching problem and presents optimal deterministic and randomized algorithms. Since then, elegant algorithmic results have been devised to accommodate more diverse environments. See the recent survey by [Meh13] and the references therein for a more comprehensive introduction.
In recent years, another online matching paradigm, online (bipartite) minimum cost matching, has emerged. Motivated by the application of ride-hailing, in this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. Upon the arrival of a request $i$, it has to be immediately and irrevocably assigned to an available server $j$, incurring a cost $\delta(i,j)$, and both request $i$ and server $j$ are removed [^1] We always assume $m\geq n$ to guarantee that every request can be served.. The objective is to minimize the total matching cost.
Due to the importance of application to ride-hailing, it is reasonable to assume that both servers and requests are located in the same metric space, and the cost incurred by matching a server and a request is measured by the distance between them. A majority of literature on online metric matching focuses on the adversarial model [PS23, Rag18, MNP06], where the locations of all servers and requests are adversarially chosen, or the weaker random order model [Rag16], which assumes that the order of adversarially chosen requests is randomly permuted. See Section 1.2 for a more comprehensive literature review in these settings.
The online matching model with adversarial online arrivals is usually deemed too pessimistic since practical instances generally exhibit a more stochastic nature. To model a practical and hence more tractable scenario, the stochastic online metric matching problem, where the requests are independently drawn from a known probability distribution, has attracted considerable attention in the past few years [GGPW19, CKKZ23, BFP23]. The most basic model assumes balanced markets, which contain an equal number of servers and requests. For the Euclidean space $[0,1]^{d}$, where $d$ is a constant not scaling with $n$, [Kan22] proposes the Hierarchical Greedy algorithm with a constant competitive ratio when all servers and requests are uniformly distributed, and [CKKZ23] achieves nearly optimal regret guarantees when servers and requests follow different distributions. [BFP23] analyzes the performance of the Greedy algorithm, whose prominence is empirically validated, and shows that Greedy is $O(1)$-competitive when servers and requests are uniformly distributed in $[0,1]$.
[AALS22] notices that the potential value of excess servers has been overlooked by the literature and considers unbalanced markets, which possess at most a linear amount of excess servers. When servers and requests are uniformly distributed in $[0,1]$, they show that the total matching cost of Greedy can be substantially improved by transitioning from balanced markets to unbalanced markets, and their results imply that Greedy is $O(\log^{3}n)$-competitive. [BFP23] subsequently pinpoints the competitive ratio of Greedy to be $O(1)$ under the same setting. Moreover, Hierarchical Greedy also admits a constant competitive ratio in unbalanced markets when servers and requests are uniformly distributed in $[0,1]^{d}$ for all $d\geq 1$ [Kan22].
In contrast to the fruitful achievements when both servers and requests are stochastic, it appears to be more challenging to establish positive results for adversarial servers and stochastic requests. Specifically, in balanced markets, [GGPW19] gives an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, which further achieves a constant competitive ratio for tree metrics. Moreover, in balanced markets with requests uniformly distributed in $[0,1]$, [BFP23] shows that Greedy is $\Theta(\log n)$-competitive and Hierarchical Greedy is $\Omega(n^{1/4})$-competitive.
1.1 Our Results
@@
As our main contribution, assuming requests are stochastic, we present a reduction from the setting with adversarial servers to that with stochastic servers: If an algorithm ${\mathcal{A}}$ exists for balanced markets with both servers and requests randomly drawn from the same distribution $\mathbb{D}$, we can use ${\mathcal{A}}$ in a black-box manner to derive another algorithm ${\mathcal{A}}^{\prime}$ for unbalanced markets with servers adversarially chosen and requests randomly drawn from $\mathbb{D}$ (Theorem 1). Moreover, if the markets are balanced, which can be viewed as a special case of unbalanced markets, ${\mathcal{A}}^{\prime}$ admits the same competitive ratio as ${\mathcal{A}}$ up to a constant factor. Hence, by utilizing the recent progress on the setting with stochastic servers and requests [Kan22, CKKZ23], and applying our reduction techniques, we then describe the following results.
@@
Table 1: Competitive ratios for the Euclidean space $[0,1]^{d}$ with adversarial servers and stochastic requests. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratio of $O(1)$ for $d=1$ is obtained by [GGPW19] and Theorem 9, which implies the two bounds on its left, and the competitive ratio of $O((\log\log\log n)^{2})$ for $d\geq 2$ is achieved by [GGPW19], which implies all $O((\log\log\log n)^{2})$ bounds on its left.
Multiplicative bound (competitive ratio).
@@
We first introduce the competitive ratio results for the Euclidean space $[0,1]^{d}$ achieved by combining our reduction with algorithms of prior work, which are summarized in Table 1. Recall that the competitive ratio of an algorithm is defined as the worst ratio between the cost incurred by the algorithm and the optimal offline cost, and can be viewed as a multiplicative guarantee. We start by showing that combining our reduction with the Hierarchical Greedy algorithm of [Kan22] leads to an $O(1)$-competitive algorithm in balanced markets for all $d\geq 1$ with adversarial servers and uniformly distributed requests (Theorem 2).
@@
Next, we describe our results in balanced markets with servers adversarially chosen and requests randomly drawn from a smooth distribution, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm ${\mathcal{A}}$ required by the reduction, we adopt the $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ ($\mathsf{SOAR}$) algorithm of [CKKZ23], which is originally used to give nearly optimal regrets in balanced markets when servers and requests follow different distributions. We show that this leads to an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 3).
@@
Finally, we present our results in unbalanced markets with adversarial servers and stochastic requests. When requests are randomly drawn from a smooth distribution, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 4). To improve the undesirable competitive ratio for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when both servers and requests are adversarially located in a doubling metric [^2] A metric space is said to be doubling if a doubling constant $M>0$ exists such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean space $[0,1]^{d}$ is a doubling metric for constant $d$., is preserved in unbalanced markets (Corollary 2).
Additive bound (regret).
@@
We turn to our regret results for the Euclidean space $[0,1]^{d}$. Since the regret of an algorithm is defined as the worst difference between the average cost incurred by the algorithm and the average optimal offline cost, it can be viewed as an additive guarantee. Recall that in balanced markets with servers and requests following different but arbitrary distributions, [CKKZ23] shows that the $\mathsf{SOAR}$ algorithm admits a regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\log n)$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the $\mathsf{SOAR}$ algorithm, we show that the same regrets are achievable even in unbalanced markets with servers adversarially chosen and requests randomly drawn from an arbitrary distribution (Theorem 5). Since balanced markets form a special case of unbalanced markets and our setting with adversarial servers is more general, the lower bounds by [CKKZ23] continue to hold in our setting, and hence, our regret guarantees are nearly optimal.
Additional results.
@@
As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics with adversarial servers and requests from balanced markets to unbalanced markets. This is achieved based on the observation that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 6). As a by-product, we similarly show that the $O(\log^{3}n)$ competitive ratio of the algorithm by [MNP06] in balanced markets with servers and requests adversarially located in an arbitrary metric is retained in unbalanced markets (Corollary 1). We remark here that all prior results for adversarial servers and requests are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.
Moreover, in Appendix B, we extend our reduction to the case where cost functions only approximately satisfy the triangle inequality. Specifically, we say that a cost function $\delta$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z$, we have $\delta(x,z)\leq\eta\cdot(\delta(x,y)+\delta(y,z))$. To showcase the generality of our approach, we show that combining our generalized reduction with the $\mathsf{SOAR}$ algorithm leads to positive results for $[0,1]^{d}$ with cost function $\left\|{\cdot}\right\|_{2}^{p}$ for some $p\geq 1$.
Finally, in Appendix C, we show that combining our reduction with the $\mathsf{SOAR}$ algorithm gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions (Theorem 9), recovering the result by [GGPW19].
@@ 1.2 Related Work
For general metrics with adversarial servers and requests, [KP93], [KMV94], and [Rag16] respectively give $(2n-1)$-competitive deterministic algorithms, which is also shown to be tight for deterministic algorithms. For randomized algorithms, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06]; it is subsequently improved by [BBGN14] to $O(\log^{2}n)$, which remains the best upper bound. In the random-order model, the deterministic algorithm of [Rag16] admits a competitive ratio of $O(\log n)$, which is optimal even for randomized algorithms, and [GK19] shows that the competitive ratio of Greedy lies between $\Omega(n^{0.22})$ and $O(n)$.
@@
For the line metric with adversarial servers and requests, [KP93] shows that the competitive ratio of Greedy is at least $12^{n}-$ 1, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] presents an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. Moving to randomized algorithms, [GL12] proposes an $O(\log n)$-competitive algorithm for the more general doubling metrics. For lower bounds, [FHK05] shows that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
@@
Various papers study other aspects of stochastic online metric matching. [TTC94] shows that Greedy admits a competitive ratio of $O(\sqrt{n})$ on the unit disk. In addition, [SNL+23] characterizes the cost of Greedy in unbalanced markets with servers and requests uniformly distributed in $[0,1]$ using a partial differential equation approach.
2 Preliminaries

### v2

@@ Abstract
We study the online metric matching problem. There are $m$ servers and $n$ requests located in a metric space, where all servers are available upfront and requests arrive one at a time. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to a free server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
When servers are adversarial and requests are independently drawn from a known distribution, we reduce the problem to a more accessible setting where all servers and requests are independently drawn from the same distribution. Applying our reduction, for the Euclidean space $[0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regrets in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of Gupta et al. (ICALP’19) for balanced markets in various regimes, and provide the first positive results for unbalanced markets.
Moreover, when the servers and requests are adversarial and a prediction of request locations is provided, for balanced markets, we present a general framework for transforming an arbitrary algorithm that does not use predictions into an algorithm that leverages predictions. The competitive ratio of the resulting algorithm is a constant with a perfect prediction and degrades smoothly as the prediction accuracy deteriorates, while preserving the worst-case guarantee.
@@ 1 Introduction
The development of centralized matching markets has emerged as a crucial research area across operations research, economics, and computer science, driven by its diverse applications in kidney exchange programs, ride-sharing services, and online advertising platforms. Market designers in these domains face the fundamental challenge of making matching decisions under future uncertainty, with the primary objective of maximizing the total incurred value. The growing significance of these applications has sparked extensive research across multiple related areas. These include online bipartite matching [KVV90, Meh13, HTW24], online resource allocation [VB21, BKK23, BLM23, BBP24, Bra24, JMZ24], and dynamic matching [AAGK17, ALG20, WXY23, Gup24, KAG24, KYY25].
A particularly compelling example of these applications can be found in the recent proliferation of ride-hailing platforms such as Didi, Uber, and Lyft. In these systems, the platform must make an immediate matching decision when a customer requests a ride - significantly delaying or rejecting the request is typically not an option. The primary cost metric in each match is the distance a driver needs to travel to reach their assigned customer. While a simple greedy approach matches each arriving customer to their nearest available driver, such myopic decisions can often lead to suboptimal outcomes. For instance, after matching a driver to their nearest customer, the next customer request might arrive in that driver’s original location, requiring a much longer pickup distance from the remaining available drivers. Therefore, the platform’s key challenge is to design matching policies that minimize the total pickup cost by considering both current matches and potential future request patterns, while maintaining the constraint of immediate matching for each request.
Inspired by the special features of the market dynamics underlying ride-hailing platforms, in this paper, we focus on the online (bipartite) minimum cost matching problem. In this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. We model the instantaneity of performing a match by assuming that each arriving request has to be immediately and irrevocably assigned to an available server. Due to the importance of application to ride-hailing, we assume that all servers and requests are located in the same metric space, and the cost incurred by matching a server and a request is measured by the distance between them. After realizing a match, both the matched server and request are removed. The objective of a central market planer is to minimize the total matching cost. We will consider both balanced and unbalanced markets, where balanced markets contain an equal number of servers and requests, and unbalanced markets possess at most a linear amount of excess servers. We will adopt competitive ratio (resp. regret) as our benchmark, which is defined as the ratio (resp. difference) between the expected average cost of matching each request incurred by the algorithm and that incurred by the optimal offline solution.
A majority of literature on online metric matching focuses on the adversarial model [PS23, Rag18, MNP06], where the locations of all servers and requests are adversarially chosen, or the weaker random-order model [Rag16], which assumes that the arrival order of adversarially chosen requests is randomly permuted. Both models, albeit being fairly general, suffer from super-constant competitive ratio lower bounds [Rag16, PS23]. Recent lines of work bypass these barriers by modeling more practical and tractable scenarios in which algorithms are equipped with additional information regarding future uncertainty. Specifically, [GGPW19, BFP23] assume that requests are randomly drawn from a known distribution, and [ACE+23b] consider the case where algorithms are provided with predictions of request locations. They achieve improved or even nearly optimal results under various choices of modeling assumptions.
1.1 Main Contribution
As our main contribution, we propose a generic approach to designing algorithms for the online metric matching problem that leverages information about future uncertainty - either through probability distributions or predictions of request locations. These algorithms result from enhancing an arbitrary algorithm ${\mathcal{A}}$ that does not use additional information. A key advantage of our approach is that this enhancement only invokes ${\mathcal{A}}$ in a black-box manner without imposing technical assumptions on ${\mathcal{A}}$.
More specifically, given predictions of request locations, for an arbitrary algorithm ${\mathcal{A}}$ that does not use predictions, we can convert ${\mathcal{A}}$ into an enhanced algorithm ${\mathcal{A}}^{\prime}$ that bases its matching decisions on both the predictions and ${\mathcal{A}}$’s strategy. The performance of ${\mathcal{A}}^{\prime}$ depends jointly on the prediction accuracy and ${\mathcal{A}}$’s performance (Theorem 1). When given distributional information instead, we generate predictions by sampling from the distribution, allowing us to apply the same approach.
We then demonstrate the effectiveness of our framework through several significant results in the stochastic and prediction settings respectively, following by our additional results for the adversarial setting.
Stochastic Setting.
@@ 1.1 Main Contribution
When servers are adversarially chosen and requests are independently drawn from a known distribution $\mathbb{D}$, we show that the aforementioned approach results in a reduction to the setting with stochastic servers and requests. Specifically, if there is an algorithm ${\mathcal{A}}$ for balanced markets with servers and requests all independently drawn from $\mathbb{D}$, we can use ${\mathcal{A}}$ in a black-box manner to derive another algorithm ${\mathcal{A}}^{\prime}$ for unbalanced markets with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ (Theorem 2). Moreover, if the markets are balanced, which can be viewed as a special case of unbalanced markets, ${\mathcal{A}}^{\prime}$ admits the same competitive ratio as ${\mathcal{A}}$ up to a constant factor. Next, we describe our results in more concrete settings obtained by further utilizing the recent progress in the setting with stochastic servers and requests [Kan22, CKKZ23].
@@
Table 1: Competitive ratios for the Euclidean space $[0,1]^{d}$ with adversarial servers and stochastic requests. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratios of $O(1)$ for $d=1$ and of $O((\log\log\log n)^{2})$ for $d\geq 2$ are obtained by [GGPW19], which imply the bounds on their left.
@@
We first introduce the competitive ratio results for the Euclidean space $[0,1]^{d}$ achieved by combining our reduction with algorithms of prior work, which are summarized in Table 1. Prior to our work, the best algorithm in balanced markets with general metrics and distributions achieves a competitive ratio of $O((\log\log\log n)^{2})$, and the same algorithm is known to be $O(1)$-competitive for tree metrics [GGPW19]. We start by showing that combining our reduction with the Hierarchical Greedy algorithm of [Kan22] leads to an $O(1)$-competitive algorithm in balanced markets for all $d\geq 1$ with adversarial servers and uniformly distributed requests (Theorem 4).
@@
Next, we describe our results in balanced markets with servers adversarially chosen and requests independently drawn from a smooth distribution, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm ${\mathcal{A}}$ required by the reduction, we adopt the Simulate-Optimize-Assign-Repeat (SOAR) algorithm of [CKKZ23], which results in an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 5).
@@
Finally, we present our results in unbalanced markets with adversarial servers and stochastic requests. When requests are independently drawn from a smooth distribution, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 6). To improve the undesirable competitive ratio for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when all servers and requests are adversarially located in a doubling metric [^1] A metric space is said to be doubling if a doubling constant $M>0$ exists such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean space $[0,1]^{d}$ is a doubling metric for constant $d$., is preserved in unbalanced markets (Corollary 3). These constitute the first known results in unbalanced markets for adversarial servers and stochastic requests.
@@
We further describe our additional results with respect to regret analysis, which are for the Euclidean space $[0,1]^{d}$. Recall that in balanced markets with servers and requests following different and arbitrary distributions, [CKKZ23] show that the SOAR algorithm admits a regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\log n)$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the SOAR algorithm, we show that the same regrets are achievable even in unbalanced markets with servers adversarially chosen and requests independently drawn from an arbitrary distribution (Theorem 7). Since balanced markets form a special case of unbalanced markets, and our setting with adversarial servers is more general, the lower bounds by [CKKZ23] continue to hold in our case, implying our regret guarantees to be nearly optimal.
Prediction Setting.
In the prediction setting, algorithms receive a set of $n$ points as a prediction at the beginning, which predicts the unknown request locations. We will focus on balanced markets with servers and requests being adversarially chosen. In particular, the prediction is not guaranteed to be correct with its error oblivious to algorithms, and algorithms are expected to be (1) consistent: the performance of algorithms is nearly optimal with a perfect prediction, and (2) robust: when the prediction admits a large error, the performance of algorithms is comparable to that of the best online algorithm that does not use predictions. In our case, the prediction error $\eta$ is defined as the minimum cost of all perfect matchings between the set of predicted request locations and the set of actual request locations.
Given an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions, we apply our generic approach mentioned at the beginning of this subsection to derive an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and achieves a competitive ratio of $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$, where $\mathrm{OPT}$ denotes the optimal offline cost (Theorem 8). To see the consistency of ${\mathcal{A}}^{\prime}$, its competitive ratio converges to a constant when the prediction error $\eta$ approaches 00. Moreover, the competitive ratio of ${\mathcal{A}}^{\prime}$ degrades smoothly as the prediction error increases, and the competitive ratio guarantee of $\alpha$ provided by ${\mathcal{A}}$ is maintained up to a constant factor regardless of the prediction error, indicating the robustness of ${\mathcal{A}}^{\prime}$. Finally, we apply the above result to several specific settings (Corollary 1).
Adversarial Setting.
@@
Now, we discuss our results for the setting in which servers and requests are adversarially chosen with no predictions provided. As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics from balanced markets to unbalanced markets. This is achieved based on the observation that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 9). As a by-product, we similarly show that the $O(\log^{3}n)$ competitive ratio of the algorithm by [MNP06] in balanced markets with servers and requests adversarially located in an arbitrary metric is retained in unbalanced markets (Corollary 2). We remark here that all prior results for adversarial servers and requests are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.
@@ 1.2 Related Work
A majority of literature for online metric matching focuses on the adversarial setting in balanced markets. For general metrics with adversarial servers and requests, [KP93], [KMV94], and [Rag16] respectively give $(2n-1)$-competitive deterministic algorithms, which is also shown to be tight. When randomization is allowed, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06]; it is subsequently improved by [BBGN14] to $O(\log^{2}n)$, which remains the best upper bound. In the random-order model, the deterministic algorithm of [Rag16] admits a competitive ratio of $O(\log n)$, which is optimal even for randomized algorithms, and [GK19] show that the competitive ratio of the Greedy algorithm lies between $\Omega(n^{0.22})$ and $O(n)$.
@@
For the line metric with adversarial servers and requests, [KP93] show that the competitive ratio of Greedy is at least $12^{n}-$ 1, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] present an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. Moving to randomized algorithms, [GL12] propose an $O(\log n)$-competitive algorithm for the more general doubling metrics. For lower bounds, [FHK05] show that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
Stochastic Online Metric Matching.
The stochastic online metric matching problem, where the requests are independently drawn from a known probability distribution, has attracted considerable attention in the past few years. We start with the references for balanced markets. For the Euclidean space $[0,1]^{d}$, where $d$ is a constant not scaling with $n$, [Kan22] proposes the Hierarchical Greedy algorithm with a constant competitive ratio when all servers and requests are uniformly distributed, and [CKKZ23] achieve nearly optimal regret guarantees when servers and requests are allowed to follow different distributions. [BFP23] show that Greedy is $O(1)$-competitive when servers and requests are uniformly distributed in $[0,1]$.
Motivated by the potential value of excess servers, [AALS22] initiate the study of unbalanced markets. When servers and requests are uniformly distributed in $[0,1]$, they show that the total matching cost of Greedy can be substantially improved by transitioning from balanced markets to unbalanced markets, and their results imply that Greedy is $O(\log^{3}n)$-competitive in unbalanced markets. [BFP23] subsequently pinpoint the competitive ratio of Greedy to be $O(1)$ under the same setting. Moreover, Hierarchical Greedy also admits a constant competitive ratio in unbalanced markets when servers and requests are uniformly distributed in $[0,1]^{d}$ for all $d\geq 1$ [Kan22].
In contrast to the fruitful achievements when both servers and requests are stochastic, much less results are known for adversarial servers and stochastic requests. In balanced markets, [GGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, which further achieves a constant competitive ratio for tree metrics. Moreover, in balanced markets with requests uniformly distributed in $[0,1]$, [BFP23] show that Greedy is $\Theta(\log n)$-competitive and Hierarchical Greedy is $\Omega(n^{1/4})$-competitive.
@@
Various papers study other aspects of stochastic online metric matching. [TTC94] show that Greedy admits a competitive ratio of $O(\sqrt{n})$ on the unit disk. In addition, [SNL+23] characterize the cost of Greedy in unbalanced markets with servers and requests uniformly distributed in $[0,1]$ using a partial differential equation approach.
Online Algorithm with Predictions.
In recent years, due to the unparalleled impact of machine learning, the paradigm of designing online algorithms using externally provided predictions is of particular interest to overcome both information-theoretic and computational barriers. This paradigm suitably models the practical scenarios where future uncertainty is predictable according to prior knowledge, where predictions can be generated by machine learning models, human experts, historical data, heuristics, or even crowdsourcing. In particular, this line of literature usually assumes the provided predictions are untrusted with their error oblivious to algorithms, and algorithms are required to be consistent and robust. Besides, the performance of algorithms usually deteriorates smoothly as a function of the prediction error when the prediction accuracy decreases.
This paradigm has been incorporated into the study of various classical problems including online metric matching [ACE+23b], metrical task systems [ACE+23b, ACE+23a], online facility location [JLL+22, ACL+21], online bipartite matching [AGKK23, JM22], and caching [GPSS22, AGKP22, LV21], among others.
It is worth of mentioning that [ACE+23b] also study online metric matching with predictions. As the primary difference, we assume that the predictions are request locations and given at the beginning, whereas in their model, each arriving request is accompanied by a prediction of the behaviors of the optimal offline algorithm until the end of the current round, i.e., all servers matched by the optimal offline algorithm for the requests arrived so far. [^2] They allow the predictions to be inconsistent, i.e., each prediction may not be a superset of previous predictions. Moreover, their prediction error is defined against an optimal offline algorithm as the summation of prediction errors of all rounds, where the prediction error of each round is defined as the minimum cost of moving the predicted matched servers to the actual servers matched by the optimal offline algorithm until the end of this round. Hence, their results are in general incomparable with ours.
We note that concurrent and independent work by [CCM25] gives a consistent (but non-robust) algorithm for online metric matching with predictions that is very similar to ours. In particular, they assume that the prediction is given as a distribution over the metric space, and their algorithm draws $n$ independent samples from this distribution as predicted request locations, whereas we directly assume that $n$ points in the metric space are given as a prediction.
2 Model Setup